# Unsupervised Density Neural Representation for CT Metal Artifact Reduction

### v1

@@ Abstract
Emerging unsupervised reconstruction techniques based on implicit neural representation (INR), such as NeRP, CoIL, and SCOPE, have shown unique capabilities in CT linear inverse imaging. In this work, we propose a novel unsupervised density neural representation (Diner) to tackle the challenging problem of CT metal artifacts when scanned objects contain metals. The drastic variation of linear attenuation coefficients (LACs) of metals over X-ray spectra leads to a nonlinear beam hardening effect (BHE) in CT measurements. Recovering CT images from metal-affected measurements therefore poses a complicated nonlinear inverse problem. Existing metal artifact reduction (MAR) techniques mostly formulate the MAR as an image inpainting task, which ignores the energy-induced BHE and produces suboptimal performance. Instead, our Diner introduces an energy-dependent polychromatic CT forward model to the INR framework, addressing the nonlinear nature of the MAR problem. Specifically, we decompose the energy-dependent LACs into energy-independent densities and energy-dependent mass attenuation coefficients (MACs) by fully considering the physical model of X-ray absorption. Using the densities as pivot variables and the MACs as known prior knowledge, the LACs can be accurately reconstructed from the raw measurements. Technically, we represent the unknown density map as an implicit function of coordinates. Combined with a novel differentiable forward model simulating the physical acquisition from the densities to the measurements, our Diner optimizes a multi-layer perception network to approximate the implicit function by minimizing predicted errors between the estimated and real measurements. Experimental results on simulated and real datasets confirm the superiority of our unsupervised Diner against popular supervised techniques in MAR performance and robustness. To the best of our knowledge, our Diner is the first unsupervised MAR method that significantly outperforms its supervised counterparts.
@@
X-ray CT, Metal Artifact Reduction, Neural Representation, Unsupervised Learning
@@ I Introduction
X-ray computed tomography (CT) is an essential biomedical imaging technique for visualizing the anatomy of observed objects. CT reconstruction aims to solve the spatial distribution of linear attenuation coefficients (LACs) of the objects to X-rays. The LACs of the biological tissue remain nearly constant over X-ray spectra, while those of metals exhibit significant variations. This physical discrepancy causes a non-negligible nonlinear beam hardening effect (BHE) in CT measurements when the objects contain metals [1]. Hence, recovering the LAC map from metal-affected measurements poses a complicated nonlinear inverse problem. Traditional linear algorithms, such as filtered back-projection (FBP) [2], often cause severe artifacts in the CT images. These artifacts could significantly reduce the reliability of clinical diagnosis and biomedical research, such as bone analysis.
@@
As an essential and longstanding issue in CT imaging, many efforts have been made to tackle metal artifacts reduction (MAR) [3]. Model-based [4, 5, 6] methods formulate the MAR as an image inpainting problem. They consider metal-corrupted signals in the raw measurements as missing data and design different schemes (e.g., linear interpolation) to fill in these missing data. However, such paradigms typically cause apparent secondary artifacts in the reconstructions since the interpolation algorithms cannot meet the geometry constraints of CT scanning. Recently, supervised deep learning (DL) approaches [7, 8, 9, 10, 11, 12, 13] dominate the MAR task, in which deep neural networks normally are trained for learning mappings from the metal-effected measurements to clear measurements or the artifacts-corrupted images to the desired high-quality images. Benefiting from powerful neural networks and data-driven priors, supervised DL methods produce state-of-the-art (SOTA) MAR performance. However, they still face two major limitations: 1) A large-scale training dataset comprising numerous paired samples is required for supervised learning. However, collecting such paired datasets is a very challenging task; 2) Supervised learning often suffers from performance drops when CT acquisition protocols (e.g., , differences in metal shapes and X-ray acquisition geometry) deviate from those in the training dataset, i.e., out-of-domain (OOD) problem, The two limitations considerably hamper the practical use of the supervised MAR models in real-world scenarios.
Implicit neural representation (INR) has recently emerged as an innovative unsupervised framework for addressing inverse problems in image reconstruction. It first represents an unknown image as an implicit function that maps spatial coordinates into image pixels. Through the integration of differentiable physical forward models (e.g., Fourier transform for MRI and Radon transform for parallel CT), the unknown image can be transformed into the measurement domain (e.g., K-space data for MRI and projection data for CT). Using a multi-layer perception (MLP) to approximate the implicit function allows the unknown image to be solved by minimizing predicted errors between the estimated and real measurements. The unsupervised learning makes the INR great practical for numerous reconstruction problems. Moreover, the MLP networks induce an inherent learning bias towards low-frequency signals [14, 15]. This bias can be viewed as an implicit regularization prior, enabling the INR to provide a stable and feasible solution for under-determined inverse problems. By simplifying the CT forward acquisition as a linear integral (i.e., Radon transformation), the INR has demonstrated great potential for various CT reconstruction tasks (e.g., sparse-view CT [16, 17, 18, 19, 20, 21, 22], limited-angle CT [20, 22], and dynamic CT [23, 24, 25]). However, existing INR-based methods fail to handle the nonlinear MAR problem since the current linear integral model supposes that LACs are monochromatic strictly (i.e., energy-independent), which leads to the nonlinear BHE being ignored.
In this paper, we propose a novel unsupervised density neural representation (Diner) to tackle the MAR problem from a nonlinear perspective. Our Diner proposes to introduce an energy-dependent polychromatic CT forward model (Eq. (3)) to the INR framework, enabling its capacity to addressee the nonlinear nature of the MAR problem. Specifically, by fully considering the physical model of X-ray absorption, we disentangle the energy-dependent LAC map into two essential components: an energy-independent density map and an energy-dependent mass attenuation coefficient (MAC) map. Using the MAC map as known prior knowledge, we then use the INR network to solve the unknown density map. Therefore, the polychromatic LAC maps at different X-ray energy levels can be accurately generated by performing the polychromatic CT forward model on the estimated density map. As reliable predictions for polychromatic LAC maps can be achieved, our Diner fundamentally addresses the energy-induced nonlinear BHE and thus facilitates effective MAR reconstructions.
@@
Technically, we represent the unknown density map as an implicit function of spatial coordinates. Then, we optimize an MLP network to learn the function from the metal-affected measurements. We make two key technical contributions to perform the challenging nonlinear optimization problem. Firstly, we propose a novel differentiable physical forward model that can accurately simulate the acquisition process from the underlying densities to the measurements. Secondly, we incorporate the proposed forward model with the INR framework, enabling its capacity to solve the nonlinear inverse problem. Our Diner is a fully unsupervised technique and thus does not require any external training data, improving its great potential for a variety of clinical applications. We evaluate the proposed Diner model on four datasets, including two simulated datasets and two real-world datasets. Empirical results demonstrate that our unsupervised Diner performs slightly better than well-known supervised MAR techniques on in-domain datasets while significantly outperforming them on out-of-domain datasets.
The main contributions of this work are as follows:
@@
  • We propose Diner, an unsupervised physical-model-driven DL approach, that can recover high-quality CT images from metal-corrupted measurements without involving external training data.
@@
  • We propose a novel formulation for the nonlinear MAR problem, that is, solving the energy-independent density map, which fundamentally addresses the energy-induced nonlinear BHE.
@@
  • We propose a novel differentiable forward model that can accurately simulate the physical acquisition process from the underlying densities to the CT measurements.
@@
  • We incorporate the proposed forward model into the INR framework, enabling its capacity to handle the nonlinear MAR inverse problems.
This study follows our previous study Polyner [26] and performs several substantial improvements. Polyner [26] tackles the MAR problem by directly representing the polychromatic LAC maps at multiple energy levels using the INR framework. This significantly alleviates the energy-induced BHE. However, the prediction of numerous (##\## of multiple energy levels) LAC maps poses a severely ill-posed under-determined inverse problem. In Polyner [26], the solution is merely constrained by an extra local smooth regularization. Compared with Diner, Polyner [26] has a significantly higher number of unknown variables, and the regularization is not sufficiently robust. Consequently, this inadequacy makes it challenging to ensure a reliable prediction of the polychromatic LAC maps. In comparison, Diner proposes a novel strategy to unravel the energy-dependent LAC map by utilizing the density map and MAC map. This approach reduces the number of unknown variables to a single-density map, substantially reducing the under-determination level. A better MAR reconstruction thus can be yielded. Moreover, this work conducts a more comprehensive experimental evaluation (including multiple types of simulated metal materials, real-world 3D cone-beam mouse data, real-world 3D helical phantom data following clinical CT acquisition protocol, extensive ablation studies, failure cases, etc.), which significantly strengthens the reliability of the proposed method.
@@ II Preliminaries
$I(\mathbf{r},E)=I_{0}(E_{0})\cdot\text{exp}\left(-\int_{\mathbf{r}}\mu(\mathbf{x},E_{0})\mathrm{d}\mathbf{x}\right),$ (1)
where $\mathbf{r}$ represents a monochromatic (i.e., single-energy) X-ray, $E_{0}$ is its energy level, $I_{0}(E)$ and $I(\mathbf{r},E)$ respectively are the numbers of photons emitted by an X-ray source and received by an X-ray detector, and $\mu(\mathbf{x},E)$ is the LAC of the object at the position $\mathbf{x}$. The LAC value serves as a measure of the capacity of the object to absorb the X-ray.
The CT measurements $p(\mathbf{r})$ are typically defined as a linear integral of the LACs $\mu$ along the X-ray $\mathbf{r}$:
$p(\mathbf{r})\triangleq-\ln\frac{I(\mathbf{r},E_{0})}{I_{0}(E_{0})}=\int_{\mathbf{r}}\mu(\mathbf{x},E_{0})\mathrm{d}\mathbf{x}.$ (2)
The CT reconstruction problem is to solve the spatial distribution of the LACs $\mu(\mathbf{x},E_{0})$ from the measurement data $p$. When the Nyquist-Shannon sampling theorem is satisfied [29], standard techniques (e.g., FBP [2]) for linear inverse problems can produce high-quality CT images.
However, most clinical CT scanners emit polychromatic X-rays (i.e., within a range of energy levels) due to physical limitations of the X-ray generator hardware [1]. Given an X-ray composed of multiple monochromatic X-rays within a range of $\mathcal{E}$, the CT measurement data can be expressed by a nonlinear polychromatic model:
$p(\mathbf{r})$ $=-\ln\frac{\int_{\mathbf{\mathcal{E}}}I(\mathbf{r},E)\mathrm{d}E}{\int_{\mathbf{\mathcal{E}}}I_{0}(E)\mathrm{d}E}$ (3)
@@
$p(\mathbf{r})=-\ln\int_{\mathcal{E}}\eta(E)\cdot\mathrm{exp}\left(-\int_{\mathbf{r}}\mu(\mathbf{x})\mathrm{d}\mathbf{x}\right)\mathrm{d}E=\int_{\mathbf{r}}\mu(\mathbf{x})\mathrm{d}\mathbf{x}.$ (5)
@@
However, the LACs of metallic materials undergo sharp variations as a function of the energy spectrum of the X-rays [1]. For an organism with metallic implants, we thus have to split its LAC map into two parts: tissue region and metal region. Mathematically, it can be expressed as:
@@
$p(\mathbf{r})=$ $\overbrace{\int_{\mathbf{r}}\left[1-\mathcal{M}(\mathbf{x})\right]\cdot\mu_{\mathrm{tissue}}(\mathbf{x})\mathrm{d}\mathbf{x}}^{\text{\small{Linear Integral}}}$ (7)
@@
Eq. (7) means that the CT measurements of an organism with metallic implants should be divided into two parts: (1) a linear integral for the tissue and (2) an energy-induced non-linear transformation for the metals. The latter is known as the nonlinear beam hardening effect (BHE) in the CT measurements. In this case, the use of traditional linear algorithms, such as FBP [2], will result in severe metal artifacts in the resulting CT images.
@@ III-A Problem Formulation
As illustrated in Eq. (7), the nonlinear BHE is the key factor for addressing the CT MAR problem. Previous MAR approaches [31, 7, 32, 4, 5] mostly assume that the LACs to be resolved are energy-independent and reconstruct them at a virtual monochromatic energy level. However, this paradigm fails to consider the energy-induced BHE, which results in the sub-optimal MAR performance.
In this work, we propose to directly reconstruct the densities of observed objects. The densities are physically energy-independent, which can fundamentally address the energy-induced BHE and thus significantly improve the MAR performance. As shown in Eq. (3), the CT measurement data $p$, however, is the nonlinear transform of the energy-dependent LACs $\mu$. It is not directly related to the densities. To establish an explicit relationship between the measurement data $p$ and densities, we innovatively introduce a standard physical model [33, 34] to decompose the LACs $\mu$ as a product of the densities $\sigma$ and energy-dependent mass attenuation coefficients (MACs) $\gamma$, which is expressed as below:
@@
For organisms with metallic implants, we then make two basic assumptions:
@@
The MACs of the biological tissue can be approximated by those of water, i.e., $\gamma(\mathbf{x},E)\triangleq\gamma_{\mathrm{water}}(E),\forall\mathbf{x}\notin\mathcal{M}$, where $\mathcal{M}$ denotes metal regions.
@@
Most common biomedical metallic implants (e.g., titanium and 304 stainless steel) are homogeneous, and their corresponding MACs can be easily accessible in some well-known open-source databases, such as NIST standard library [37] and XCOM program [30].
According to assumptions 1 and 2, by substituting Eq. (8) into Eq. (6), we define a novel LACs decomposition model for the MAR task as below:
@@
where $\gamma_{\mathrm{water}}$ and $\gamma_{\mathrm{metal}}$ respectively denote the known MACs of the water and metals. Eq. (9) implies that the energy-dependent LACs $\mu(\mathbf{x},E)$ can be reconstructed accurately by solving the energy-independent densities $\sigma(\mathbf{x})$. Hence, the CT MAR problem is formulated as a reconstruction problem of the densities of the observed objects.
Figure 2: Overview of the proposed Diner model. For an X-ray $\mathbf{r}$, we first sample a set of coordinates $\mathbf{x}$ at a fixed interval $\Delta\mathbf{x}$. Then, we feed these coordinates $\mathbf{x}$ into an MLP network $g_{\phi}$ to predict the densities $\sigma(\mathbf{x})$ of the observed object at these positions. Furthermore, we use the proposed nonlinear forward model $\mathcal{T}$ (Eq. (12)) to transform these MLP-predicted densities $\sigma(\mathbf{x}),\forall\mathbf{x}\in\mathbf{r}$ into measurement data $\hat{p}(\mathbf{r})$. Finally, we minimize the predicted errors $\mathcal{L}_{\text{DC}}$ between the estimated $\hat{p}(\mathbf{r})$ and raw $p(\mathbf{r})$ measurements to optimize the trainable weights of the MLP network $g_{\phi}$ for learning the density neural representation.
III-B Density Neural Representation

### v2

# Solving Energy-Independent Density for CT Metal Artifact Reduction via Neural Representation

@@ Abstract
X-ray CT often suffers from shadowing and streaking artifacts in the presence of metallic materials, which severely degrade imaging quality. Physically, the linear attenuation coefficients (LACs) of metals vary significantly with X-ray energy, causing a nonlinear beam hardening effect (BHE) in CT measurements. Reconstructing CT images from metal-corrupted measurements consequently becomes a challenging nonlinear inverse problem. Existing state-of-the-art (SOTA) metal artifact reduction (MAR) algorithms rely on supervised learning with numerous paired CT samples. While promising, these supervised methods often assume that the unknown LACs are energy-independent, ignoring the energy-induced BHE, which results in limited generalization. Moreover, the requirement for large datasets also limits their applications in real-world scenarios. In this work, we propose Density neural representation (Diner), a novel unsupervised MAR method. Our key innovation lies in formulating MAR as an energy-independent density reconstruction problem that strictly adheres to the photon-tissue absorption physical model. This model is inherently nonlinear and complex, making it a rarely considered approach in inverse imaging problems. By introducing the water-equivalent tissues approximation and a new polychromatic model to characterize the nonlinear CT acquisition process, we directly learn the neural representation of the density map from raw measurements without using external training data. This energy-independent density reconstruction framework fundamentally resolves the nonlinear BHE, enabling superior MAR performance across a wide range of scanning scenarios. Extensive experiments on both simulated and real-world datasets demonstrate the superiority of our unsupervised Diner over popular supervised methods in terms of MAR performance and robustness. To the best of our knowledge, Diner is the first unsupervised MAR method to outperform its supervised counterparts.
@@
X-ray CT, Metal Artifact Reduction, Neural Representation, NeRF, Unsupervised Learning
@@ I Introduction
X-ray computed tomography (CT) is a vital biomedical imaging technique for visualizing the anatomy of objects. CT imaging aims to determine the spatial distribution of the linear attenuation coefficients (LACs) of objects to X-rays. While the LACs of biological tissues remain nearly constant across X-ray energies, those of metals exhibit significant variations. This physical discrepancy introduces a substantial nonlinear beam hardening effect (BHE) in CT measurements when metals are present [1]. Consequently, recovering the LAC map from metal-affected measurements becomes a complex nonlinear inverse problem. Traditional linear algorithms, such as FBP [2], often produce severe shadowing and streaking artifacts in CT images. These artifacts can significantly compromise the reliability of clinical diagnoses and biomedical research, including applications such as bone analysis.
@@
Metal artifact reduction (MAR) has long been a critical and enduring challenge in X-ray CT [3]. Model-based methods [4, 5] approach MAR by formulating it as an image inpainting problem. These methods treat metal-corrupted signals in the measurements as missing data and employ various schemes (e.g., linear interpolation) to fill in the gaps. However, such frameworks often introduce noticeable secondary artifacts in the reconstructed images, as the interpolation algorithms fail to satisfy the geometric constraints of CT scanning. Recently, supervised deep learning (DL) models [6, 7, 8, 9, 10, 11, 12] have dominated the MAR problem. In these methods, deep neural networks are trained to map metal-affected measurements to clean measurements or artifact-corrupted images to artifact-free images. By leveraging the capabilities of neural networks and data-driven priors, supervised DL methods achieve state-of-the-art (SOTA) MAR performance. However, they face two major limitations: 1) Supervised learning requires a large-scale training dataset with numerous paired samples, which is extremely challenging to collect; 2) These methods are prone to performance degradation when CT acquisition protocols (e.g., variations in metal shapes and X-ray acquisition geometry) deviate from those used in the training dataset, leading to the out-of-domain (OOD) problem. These limitations significantly hinder the practical application of supervised MAR models in real-world scenarios.
Implicit neural representation (INR) has recently emerged as a revolutionary unsupervised DL framework for various CT reconstruction tasks, such as sparse-view CT [13, 14, 15, 16, 17, 18], limited-angle CT [19, 18, 20], and dynamic CT [21, 22, 23]. INR represents an unknown CT image as a continuous function of spatial positions. By incorporating a differentiable linear integral model (i.e., Radon transform), it optimizes a multi-layer perceptron (MLP) to approximate the function by minimizing the errors in the measurements. Due to the inherent learning bias of MLPs toward low-frequency image patterns [24, 25], INR can generate high-quality CT images under highly ill-posed conditions. Recently, Wu et al.[26] proposed Polyner, an INR-based MAR method. Its core innovation is the introduction of a polychromatic physical model to characterize the nonlinear CT acquisition process. By using the polychromatic forward model, Polyner[26] optimizes INR to reconstruct polychromatic LACs at multiple X-ray energies from raw measurements. This approach significantly mitigates the energy-induced BHE and achieves promising results. However, predicting multiple LAC maps poses a highly ill-posed inverse problem. To address this, Polyner [26] requires a local energy smoothness regularizer to constrain the solution space, which significantly limits its MAR performance and robustness.
In this work, we propose Density neural representation (Diner), a new unsupervised method that formulates MAR as an energy-independent density reconstruction problem. This formulation fundamentally addresses the nonlinear BHE, resulting in superior MAR reconstructions. Conceptually, we decompose energy-dependent LACs into the product of energy-dependent mass attenuation coefficients (MACs) and energy-independent densities using a standard physical model [27, 28]. By further employing the water-equivalent tissue approximation [29, 30] (i.e., the MACs of biological tissues are approximately equal to those of water), we can leverage prior knowledge of the known MACs of water and metal. Therefore, our objective is to directly reconstruct densities from metal-affected measurements. Compared to Polyner [26], which estimates multiple polychromatic LAC maps, our Diner substantially reduces the ill-posed nature of the inverse problem by reconstructing a single energy-independent density map, achieving enhanced reconstructions.
@@
To achieve stable optimization, we adopt INR as the core framework. Specifically, we represent the unknown density map as an implicit function of spatial coordinates. Subsequently, we introduce a novel differentiable polychromatic forward model that converts the densities into CT measurements while enabling gradient back-propagation. This allows us to optimize an MLP network to learn the function by minimizing the error between the estimated and actual measurements. Our Diner is a fully unsupervised technique, eliminating the need for external training data and thus enhancing its potential for diverse clinical applications. We evaluate Diner on four datasets, including two simulated and two real-world datasets. Empirical results demonstrate that our unsupervised Diner slightly outperforms well-known supervised MAR techniques on in-domain datasets and significantly surpasses them on out-of-domain datasets. The main contributions of this work are as follows:
@@
  • We propose Diner, an unsupervised, physical-model-driven MAR approach capable of recovering high-quality CT images from metal-corrupted measurements without requiring external training data.
@@
  • We introduce a novel formulation for the nonlinear MAR problem by solving the energy-independent density map, which fundamentally addresses the nonlinear, energy-induced BHE.
@@
  • We present a differentiable physical forward model that accurately simulates the nonlinear physical acquisition process from densities to CT measurements.
@@
  • We integrate the proposed physical forward model into the INR framework, enabling it to effectively address nonlinear MAR problems.
@@ II Preliminaries
$I(\mathbf{r},E)=I_{0}(E)\cdot\text{exp}\left(-\int_{\mathbf{r}}\mu(\mathbf{x},E)\mathrm{d}\mathbf{x}\right),$ (1)
where $\mathbf{r}$ represents a monochromatic (i.e., single-energy) X-ray, $E$ is its energy level, $I_{0}(E)$ and $I(\mathbf{r},E)$ respectively are the numbers of photons emitted by an X-ray source and received by an X-ray detector, and $\mu(\mathbf{x},E)$ is the LAC of the object at the position $\mathbf{x}$. The LAC value serves as a measure of the capacity of the object to absorb the X-ray.
The CT measurements $\rho(\mathbf{r})$ are typically defined as a linear integral of the LACs $\mu$ along the X-ray $\mathbf{r}$:
$\rho(\mathbf{r})\triangleq-\ln\frac{I(\mathbf{r},E)}{I_{0}(E)}=\int_{\mathbf{r}}\mu(\mathbf{x},E)\mathrm{d}\mathbf{x}.$ (2)
The CT reconstruction problem is to solve the spatial distribution of the LACs $\mu(\mathbf{x},E)$ from the measurement data $\rho(\mathbf{r})$. When the Nyquist-Shannon sampling theorem is satisfied [33], standard techniques (e.g., FBP [2]) for linear inverse problems can produce high-quality CT images.
However, most clinical CT scanners emit polychromatic X-rays (i.e., within a range of energy levels) due to physical limitations of the X-ray source hardware [1]. Given an X-ray composed of multiple monochromatic X-rays within a range of $\mathcal{E}$, the CT measurement data can be expressed by a nonlinear polychromatic model:
$\rho(\mathbf{r})$ $=-\ln\frac{\int_{\mathbf{\mathcal{E}}}I(\mathbf{r},E)\mathrm{d}E}{\int_{\mathbf{\mathcal{E}}}I_{0}(E)\mathrm{d}E}$ (3)
@@
$\rho(\mathbf{r})=-\ln\int_{\mathcal{E}}\eta(E)\cdot\mathrm{exp}\left(-\int_{\mathbf{r}}\mu(\mathbf{x})\mathrm{d}\mathbf{x}\right)\mathrm{d}E=\int_{\mathbf{r}}\mu(\mathbf{x})\mathrm{d}\mathbf{x}.$ (5)
@@
However, the LACs of metallic materials undergo sharp variations as a function of the X-ray energies [1]. For an organism with metallic implants, we thus have to split its LAC map into two parts: tissue region and metal region. Mathematically, it can be expressed as:
@@
$\rho(\mathbf{r})=$ $\overbrace{\int_{\mathbf{r}}\left[1-\mathcal{M}(\mathbf{x})\right]\cdot\mu_{\mathrm{tissue}}(\mathbf{x})\mathrm{d}\mathbf{x}}^{\text{\small{Linear Integral}}}$ (7)
@@
Eq. (7) means that the CT measurements of an organism containing metallic materials should be divided into two parts: (1) a linear integral for the tissue and (2) an energy-induced non-linear transform for the metals. The latter results in a nonlinear beam hardening effect (BHE) in the CT measurements. In this case, the use of traditional linear FBP [2], will result in severe metal artifacts in CT images.
@@ III-A Problem Formulation
As illustrated in Eq. 7, the nonlinear BHE is the key factor for addressing the CT MAR problem. Previous MAR approaches [35, 6, 36, 8, 9, 4] mostly assume that the LACs to be resolved are energy-independent and reconstruct them at a virtual monochromatic X-ray energy level. However, this framework fails to capture the energy-induced BHE, which results in suboptimal MAR performance.
In this work, we propose directly reconstructing the densities $\sigma$ of observed objects. Since the densities $\sigma$ are physically energy-independent, this approaches can fundamentally address the energy-induced BHE and significantly improve MAR performance. However, as shown in Eq. 3, the CT measurement data $\rho$ is a nonlinear transform of the energy-dependent LACs $\mu$, and is not directly related to the densities $\sigma$. To establish an explicit relationship between the measurement data $\rho$ and densities $\sigma$, we introduce a standard physical model [27, 28] to decompose the LACs $\mu$ as a product of the energy-independent densities $\sigma$ and energy-dependent mass attenuation coefficients (MACs) $\gamma$, as below:
@@
For a organism containing metallic implants, we leverage two basic and reasonable assumptions:
@@
(Water-Equivalent Tissues Approximation [29, 30]) The MACs of biological tissues are approximately replaced by those of water, i.e., $\gamma(\mathbf{x},E)\triangleq\gamma_{\mathrm{water}}(E),\forall\mathbf{x}\notin\mathcal{M}$, where $\mathcal{M}$ denotes metal regions.
@@
Most common biological and clinical metallic implants (e.g., titanium and 304 stainless steel) are homogeneous, and their corresponding MACs can be easily accessible in some well-known open-source databases, such as NIST standard library [37] and XCOM program [34].
According to assumptions 1 and 2, by substituting Eq. 8 into Eq. 6, we define a novel LAC decomposition model for the CT MAR problem as below:
@@
where $\gamma_{\mathrm{water}}$ and $\gamma_{\mathrm{metal}}$ respectively denote the known MACs of the water and metals. Eq. 9 implies that the energy-dependent LACs $\mu(\mathbf{x},E)$ can be reconstructed accurately by estimating the energy-independent densities $\sigma(\mathbf{x})$. Therefore, the CT MAR is formulated as an energy-independent density reconstruction problem (i.e., $\rho\rightarrow\sigma$).
Figure 2: Overview of the proposed Diner model. For an X-ray $\mathbf{r}$, we first sample a set of coordinates $\mathbf{x}$ at a fixed interval $\Delta\mathbf{x}$. Then, we feed these coordinates $\mathbf{x}$ into an MLP network $g_{\phi}$ to predict the densities $\sigma(\mathbf{x})$ of the observed object at these positions. Furthermore, we use the proposed nonlinear forward model $\mathcal{T}$ (Eq. 12) to transform these MLP-predicted densities $\sigma(\mathbf{x}),\forall\mathbf{x}\in\mathbf{r}$ into measurement data $\hat{\rho}(\mathbf{r})$. Specifically, these densities $\sigma(\mathbf{x})$ are first converted into polychromatic LACs $\{\mu(\mathbf{x},E_{i})\}_{i=1}^{N}$ using our LAC decomposition model (Eq. 9). Then, the measurement $\hat{p}(\mathbf{x})$ is generated from these LACs using the polychromatic model (Eq. 11). Finally, we minimize the predicted errors $\mathcal{L}_{\text{DC}}$ between the estimated $\hat{\rho}(\mathbf{r})$ and raw $\rho(\mathbf{r})$ measurements to optimize the trainable weights of the MLP network $g_{\phi}$ for learning the density neural representation.
III-B Solving Density via Neural Representation